OpenStudy (anonymous):

At what height would the gravitational potential energy have a value of 73.7 joules for a 15 kilogram object? Note that acceleration due to gravity is 9.8 meters/second2. A. 0.5 meters B. 1.2 meters C. 2.0 meters D. 3.1 meters E. 3.4 meters

OpenStudy (callisto):

PE = mgh PE = potential energy (in joules) m = mass of object (in kg) g = acc. due to gravity (9.8ms^-2) h = height (in m)

OpenStudy (anonymous):

PE=mass*g*height 73.7=15*9.8*? ?=.501